4. A 2.5 litre sample of a gas is heated from 250 C to 550C at constant pressure. Compute the volume at the higher temperature?


1
Expert's answer
2022-05-10T07:44:04-0400

V1 / T2 = V2 / T2

2.5 / 250 = V2 / 550

V2 = 5.5 L